A firm is trying to decide which of two devices to install to reduce costs. Both devices have useful
lives of 5 years and no salvage value. Device A costs $1000 and can be expected to result in $300
savings annually. Device B costs $1350 and will provide cost savings of $300 the first year, but
savings will increase by $50 annually, making the second-year savings $350, the third-year savings
$400, and so forth. With interest at 7%, which device should the firm purchase?
1
Expert's answer
2021-10-05T12:15:03-0400

A. PW of cost = $1000


PW of benefits"= 300(P/A, 6\%,5) = 300x4.212 = 1263.6"


B/C ratio = 1.263


B. PW of cost = $1000


"PW of benefits = 500(P/A, 6\%, 5) - 50(P/G, 6\%,5) = 500x4.212 - 50 x 7.934 = 1709.3"


The net flow is maximum for device B after 5 years therefore the firm should purchase Device B